# Robust Spatial Confounding Adjustment via Basis Voting

### v1

@@ Abstract
Estimating causal effects of spatially structured exposures is complicated by unmeasured spatial confounders, which undermine identifiability in spatial linear regression models unless structural assumptions are imposed. We develop a general framework for causal effect estimation that relaxes the commonly assumed requirement that exposures contain higher-frequency variation than confounders. We propose basis voting, a plurality-rule estimator — novel in the spatial literature — that consistently identifies causal effects only under the assumption that, in a spatial basis expansion of the exposure and confounder, there exist several basis functions in the support of the exposure but not the confounder. This assumption generalizes existing assumptions of differential basis support used for identification of the causal effect under spatial confounding, and does not require prior knowledge of which basis functions satisfy this support condition. We also show that the standard projection-based estimator used in other methods relying on differential support is inefficient, and provide a more efficient novel estimator. Extensive simulations and a real-world application demonstrate that our approach reliably recovers unbiased causal estimates whenever exposure and confounder signals are separable on a plurality of basis functions. Importantly, by not relying on higher-frequency variation, our method remains applicable to settings where exposures are smooth spatial functions, such as distance to pollution sources or major roadways, common in environmental studies.
@@ 1 Introduction
Estimating the true association between spatially structured exposures and outcomes is a fundamental problem in spatial statistics and environmental epidemiology. However, this task is challenging in the presence of unmeasured spatial variables that simultaneously influence both the exposure and the outcome. When such variables are not accounted for, the (conditional) effect size in standard regression models applied to point-referenced spatial data is unidentifiable, leading to biased estimates. This phenomenon is referred to as spatial confounding, and it arises frequently in real-world applications where spatial structure is intrinsic to both the predictors (exposures) and to unmeasured background processes.
@@
An early formal recognition of this issue can be found in Clayton et al. (1993), which demonstrated how including spatially varying error terms in the model changes covariate effect estimates, but this change was favored over the unadjusted estimate which is biased under confounding. Subsequent works have had differing opinions about this (Woodard et al. 1999, Reich et al. 2006, Wakefield 2007, Hodges & Reich 2010, Schnell & Papadogeorgou 2020, Khan & Calder 2022, Zimmerman & Ver Hoef 2022) with arguments put forth both for and against including spatially correlated errors in the model. A particularly influential contribution by Paciorek (2010) quantified this bias in generalized least squares (GLS) settings, relating it to the relative smoothness of the exposure and the unmeasured confounder. This work also underscored the identifiability challenges of separating the exposure effect from the influence of spatial confounding, unless certain restrictions or structural assumptions are imposed. Extensive theoretical and empirical results in Khan & Berrett (2023) largely agree with the conclusions of Paciorek (2010). Related work in spatial prediction frameworks, such as that of Page et al. (2017), arrives at similar conclusions. Gilbert et al. (2025) proved that the GLS estimator is consistent under varied forms of spatial confounding as long as the exposure has some non-spatial variation. Dupont et al. (2022) proves consistency of their spatial+ method under a similar assumption of Gaussian noise in the exposure. Gilbert et al. (2021) grounded the spatial confounding problem in the framework of causal inference, going beyond the linear regression case. Their sufficient conditions for identifiability of a causal effect under confounding also assumed some non-spatial variation in the exposure.
Things get more complicated if the exposure is only a function of space with no non-spatial component. Such examples are common in many environmental studies, e.g., studying the effect of distance to fracking mines on health (Casey et al. 2016, Currie et al. 2017) or of distances to roadways on air-quality (Karner et al. 2010). Brugge et al. (2013) explicitly raises concerns about the inability to adjust for co-smooth spatial confounders. Estimators like GLS and spatial+ whose consistency is guaranteed in the presence of noise in the exposure, may no longer work in this smooth-exposure setting. In fact, it has been proven that when exposures are too smooth relative to the confounder, the exposure effect cannot be identified even in a linear regression setting (Bolin & Wallin 2024, Datta & Stein 2025).
In the setting of smooth exposures, most methods rely on the assumption that confounders vary mainly at broad spatial scales, while exposures may display more localized, high-frequency spatial patterns. Extensive experiments Paciorek (2010) and Khan & Berrett (2023) have shown that such a situation is favorable for recovery of the exposure effect as the exposure becomes unconfounded at higher frequencies. This, along with the assumption of a linear relationship between the outcome and the exposure, common in spatial setup, solves the problem of spatial confounding. Methods that build on this idea include both parametric and semi-parametric approaches, such as those proposed in Thaden & Kneib (2018), Keller & Szpiro (2020), Dupont et al. (2022), and Guan et al. (2023).
@@
When the unmeasured confounder has the same or finer spatial variation than the exposure, existing techniques generally fail, and standard spatial estimators can perform worse than even the unadjusted estimator as seen in Paciorek (2010), Khan & Berrett (2023). This can be viewed through the lens of causal inference as a violation of the positivity assumption, which requires that all levels of exposure be observed across levels of confounders. Nonetheless, there are important cases where this assumption does not hold. As mentioned earlier, spatial contours of exposures like distances to industrial sites, greenery, or major roadways can be quite smooth, precluding the use of techniques relying on presence of high-frequency variation unique to the exposure. In fact, the notion of smoothness defined by the ordering of basis functions may not generalize to two or higher-dimensional spatial domains. For example, using kronecker product of basis functions of one-dimensional basis, creates a two-dimensional basis with two indices dictating smoothness (one along each coordinate). Approaches like those of Keller & Szpiro (2020) which model the confounder using basis functions of low-indices do not easily generalize for such basis sets.
@@
In this work, we develop a general framework for effect estimation in the spatial linear regression model that adjusts for unmeasured spatial confounders, without assuming that the confounders are smoother than the exposure. Instead, we introduce a milder plurality rule assumption, inspired by ideas in the robust instrumental variable (IV) literature. The central assumption is that, in a representation of the exposure $X$ and the confounder $U$ along a set of orthonormal basis functions, the set of ratios of the projections of $U$ and $X$ on each of the basis functions has a larger cluster around zero than around any other value. This assumption subsumes the rougher exposure assumption as a special case. Under plurality, we show that a set of candidate estimators of the regression effect, one based on each basis function, will have the mode around the true coefficient value. We prove that this enables consistent estimation of the effect forgoing the requirement of prior knowledge of which basis functions support only the exposure. We call the method basis voting, as each basis yields a candidate estimate, and the most popular candidate (mode) is the final estimate.
@@
A second contribution is improving the choice of the candidate estimator. Plurality rule based deconfounding has recently been considered for time-series data (Schur & Peters 2024), and the candidate estimator used is the ratio of the projections of the outcome and the exposure along a basis function. We show that this estimator is inefficient, and present a better candidate estimator with lower asymptotic variance and finite-sample unbiasedness. We provide consistency and asymptotic normality of a plurality of the candidate estimators and consistency of the aggregate mode estimator based on the candidates. We empirically demonstrate robustness to the choice of basis functions used for analyzing the data.
@@ 2 Problem Setup and Assumptions
Let $\mathcal{D}\subseteq\mathbb{R}^{d}$ denote a compact spatial domain, and let $\bm{S}_{n}=\{S_{1},\dots,S_{n}\}$ represent $n$ spatial locations sampled from a distribution with density $f_{S}(\cdot)$ on $\mathcal{D}$. At each location $S_{i}$, we observe an exposure value $X_{i}\coloneqq X(S_{i})$ and an outcome value $Y_{i}\coloneqq Y(S_{i})$. We denote the exposure and outcome vectors by $\bm{X}_{n}=(X_{1},\dots,X_{n})^{\top}$ and $\bm{Y}_{n}=(Y_{1},\dots,Y_{n})^{\top}$, respectively. In addition to $X$ and $Y$, we assume there exists an unmeasured confounder $U$ defined at each location, $U_{i}\coloneqq U(S_{i})$, represented by the vector $\bm{U}_{n}=(U_{1},\dots,U_{n})^{\top}$, which is correlated with the exposure and affects the outcome. Both the exposure and the unmeasured confounder are assumed to be univariate.
In keeping with the spatial statistics literature, we assume throughout a linear data generating process (DGP) for the outcome at a generic spatial location $s\in\mathcal{D}$:
@@
where $\beta$ is the exposure effect of interest, $\epsilon(s)$ is an independent mean-zero error term, and $U(s)$ is collinear with $X(s)$. If the confounder $U$ were measured, then it would be straightforward to estimate the exposure effect $\beta$, simply by regressing $Y$ in $X$ and $U$ and looking at the estimated slope for $X$. As $U$ is not measured, we cannot directly adjust for it in the model. In this paper, we propose novel assumptions under which we can estimate the true causal effect $\beta$ in the presence of the unmeasured confounding by $U$.
We will be assuming that both $X$ and $U$ are deterministic (fixed) functions of space, in $\mathcal{L}^{4}(\mathcal{D},f_{S})$, the Hilbert space of all fourth-power integrable functions with respect to the sampling measure $f_{S}$ with the inner product induced by it denoted by $\langle g,h\rangle_{f_{S}}\coloneqq\int_{\mathcal{D}}g(s)h(s)f_{S}(s)ds$ and the corresponding norm $\|\cdot\|_{f_{S}}$. Let $\mathcal{H}=\{h_{1},h_{2},\cdots\}$ denote an infinite-dimensional orthonormal basis of $\mathcal{L}^{4}(\mathcal{D},f_{S})$. Unlike prior approaches, our method does not require an assumption that the basis functions $h_{1},h_{2},\cdots$ are ordered by their degree of smoothness or roughness, with $X$ having a longer expansion than $U$.
Let $X(s)=\sum_{j=1}^{\infty}\alpha_{x}^{(j)}h_{j}(s)$ and $U(s)=\sum_{j=1}^{\infty}\alpha_{u}^{(j)}h_{j}(s)$ with finite $L^{2}$ norms: $\mathbb{E}[X(S)]^{2}=\sum_{j}\left|\alpha_{x}^{(j)}\right|^{2}<\infty$ and $\mathbb{E}[U(S)]^{2}=\sum_{j}\left|\alpha_{u}^{(j)}\right|^{2}<\infty$. By orthogonality, the projections of $X$ on each of the basis functions are given by the coefficients $\alpha_{x}^{(j)}=\langle X,h_{j}\rangle_{f_{S}}$ and similarly for $\alpha_{u}^{(j)}$. Because $X$ and $U$ are square-integrable, for any $\epsilon>0$, there exists $d\coloneqq d(\epsilon)\in\mathbb{N}$ large enough such that $\max\left\{\|X_{>d}\|_{f_{S}}^{2},\|U_{>d}\|_{f_{S}}^{2}\right\}<\epsilon$, where $\|f_{>d}\|_{f_{S}}$ is the norm of the residual of the function $f$ after projecting it on the first $d$ basis functions. In other words, the expansion of $X$ and $U$ with respect to $h_{1},h_{2},\cdots,h_{d}$ is a good approximation $(X,U)$.
Without loss of generality, we partition $\mathcal{H}$ into three disjoint sets $\mathcal{H}=\mathcal{H}_{UX}\sqcup\mathcal{H}_{X}\sqcup\mathcal{H}_{U}$ where $\mathcal{H}_{UX}$ is the set of basis functions shared by both $U$ and $X$, $\mathcal{H}_{X}$ is the set unique to $X$ and similarly for $\mathcal{H}_{U}$. Figure 1 depicts our DGP, with and without basis expansion.
@@
(a) Standard confounding setting with bidirectional link between $U$ and $X$
$\mathcal{H}$$\mathcal{H}_{X}$$\mathcal{H}_{UX}$$\mathcal{H}_{U}$UUXXYY
(b) Expanded setting with basis components (shaded grey)
@@
Assumption 1 (Direction of No Confounding).
Let $\bm{\alpha}_{x}=(\alpha_{x}^{(1)},\alpha_{x}^{(2)},\dots)$ and $\bm{\alpha}_{u}=(\alpha_{u}^{(1)},\alpha_{u}^{(2)},\dots)$ be the sequence of coefficients of $X$ and $U$ for the basis function set $\mathcal{H}$ respectively, then there exists a non-empty subset $O_{X}\subseteq\{1,\dots,d\}$ such that for all $j\in O_{X}$, we have $\alpha_{u}^{(j)}=0$ and $\alpha_{x}^{(j)}\neq 0$ where $d\in\mathbb{N}$ with $\max\left\{\|X_{>d}\|_{f_{S}}^{2},\|U_{>d}\|_{f_{S}}^{2}\right\}<\epsilon$ for some small $\epsilon>0$.
If the index $j$ of even a single basis function $h_{j}\in O_{X}$ were known, $h_{j}$ would suffice to identify $\beta$. This is because the effect of $h_{j}$ itself on $Y$ is unconfounded, and under linearity, its effect is equal to $\beta$. In practice, however, this will never be the case as $O_{X}$ is unknown. Previous methods, e.g., Keller & Szpiro (2020), have instead relied on the assumption that the highest frequency basis functions are unconfounded, i.e., essentially they assume that the basis functions are ordered in terms of increasing roughness or frequency $h_{1}\lessapprox h_{2}\lessapprox\cdots\lessapprox h_{d}$, and there exists some $t\leq d\in\mathbb{N}$ such that $\{t+1,\cdots,d\}\in O_{X}$. So, Assumption 1 is strictly weaker as, rather than requiring unconfounded high-frequency components, we only assume the existence of a non-trivial set $\{h_{j}:j\in O_{X}\}$. This formulation accommodates cases where the unconfounded variation in $X$ lies at low or intermediate spatial scales. Figure 2 shows two such scenarios where the roughness ordering of $X$ and $U$ is different and still can be framed by our formulation. This also broadens the scope for the choice of basis functions used for analysis, as we can now use basis functions that might not have a canonical ordering of roughness, like basis functions on multivariate domains $\mathbb{R}^{d}$ generated as a Kronecker product of bases on $\mathbb{R}$.
@@
A helpful way to interpret the subset of basis functions $\{h_{j}:j\in O_{X}\}$ is through the lens of instrumental variable (IV) methods in causal inference. In classical IV frameworks, one seeks an instrument $Z$ that satisfies two key properties: (i) it is correlated with the exposure $X$, and (ii) it is independent of the unmeasured confounder $U$ and the outcome $Y$, except through its effect on $X$. When such a valid instrument exists, projecting both $X$ and $Y$ onto the space spanned by $Z$ and regressing the projected outcome on the projected exposure yields a consistent estimate of the causal effect. In our spatial basis framework, the basis functions $\{h_{j}\}_{j=1}^{d}$ can be thought of as candidate instruments. Under Assumption 1, those $h_{j}$ with $j\in O_{X}$ act as valid instruments, in the sense that they influence $X$ but not $U$, and hence satisfy the classical exclusion restriction. The remaining basis functions $h_{j}$ for $j\notin O_{X}$ are invalid instruments, since they affect both $X$ and $U$, and their use would introduce bias into causal effect estimation.
To illustrate this connection, suppose we knew the set $O_{X}$. Then, we could project $X$ and $Y$ onto the span of $\{h_{j}:j\in O_{X}\}$ denoting these projections as $P_{O_{X}}X$ and $P_{O_{X}}Y$ and fit a simple linear regression:
$\hat{\beta}=\frac{\langle P_{O_{X}}X,P_{O_{X}}Y\rangle_{2}}{\langle P_{O_{X}}X,P_{O_{X}}X\rangle_{2}}$
which yields a consistent estimator of the true effect $\beta$. This procedure is directly analogous to two-stage least squares (2SLS), where the projections correspond to first-stage instrumented values. This IV-based perspective is consistent with the framework of Woodward et al. (2024), who show how a variety of existing approaches to unmeasured spatial confounding can be unified under a common IV interpretation.
@@
The challenge, of course, is that $O_{X}$ is unknown as we no longer assume that the unconfounded variation in $X$ is at the end of the frequency spectrum. Since we do not observe $U$, we cannot directly verify which basis functions satisfy the exclusion restriction. In the classical IV literature, several recent approaches have been developed to address this problem using methods such as Lasso-based IV selection, robust IV estimators, and plurality-rule estimators (e.g., Kang et al. (2016), Windmeijer et al. (2019)). In particular, plurality-based estimators operate under the assumption that the largest subset of candidate instruments satisfying a particular identifying condition (e.g., validity or consistency of $\hat{\beta}$) corresponds to the correct causal signal. A similar idea has been explored in the time-series setting with unmeasured temporal confounders by Schur & Peters (2024), who project both $X$ and $U$ onto a collection of basis functions and define aggregate estimators from these projections based on a majority-type rule (proportion of invalid candidates is less than 50%), which is a special case of the plurality rule. They frame this problem as an adversarial outlier problem and use robust regression techniques to estimate the effect size, based on a user-specified upper bound of the set of invalid candidates, which is assumed to be decreasing with sample size.
@@
Inspired by these ideas, we propose a plurality-rule estimator which is defined as the mode of a set of candidate estimators where each candidate is constructed based on each of the spatial basis functions $\{h_{j}\}_{j=1}^{d}$. This approach allows us to recover $\beta$ even without knowledge of $O_{X}$, under a plurality assumption that enough of the candidate basis functions are valid (i.e., contribute to $X$ but not $U$). In the next section, we formally define the plurality-rule assumption and our estimator, present its theoretical guarantees, and describe its advantages in comparison to existing approaches.

### v2

@@ Abstract
Estimating effects of spatially structured exposures is complicated by unmeasured spatial confounders, which undermine identifiability in spatial linear regression models unless structural assumptions are imposed. We develop a general framework for effect estimation in spatial regression models that relaxes the commonly assumed requirement that exposures contain higher-frequency variation than confounders. We propose basis voting, a plurality-rule estimator — novel in the spatial literature — that consistently identifies causal effects only under the assumption that, in a spatial basis expansion of the exposure and confounder, there exist several basis functions in the support of the exposure but not the confounder. This assumption generalizes existing assumptions of differential basis support used for identification of the causal effect under spatial confounding, and does not require prior knowledge of which basis functions satisfy this support condition. We design this estimator as the mode of several candidate estimators each computed based on a single working basis function. We also show that the standard projection-based candidate estimator typically used in other plurality-rule based methods is inefficient, and provide a more efficient novel candidate. Extensive simulations and a real-world application demonstrate that our approach reliably recovers unbiased causal estimates whenever exposure and confounder signals are separable on a plurality of basis functions. By not relying on higher-frequency variation, our method remains applicable to settings where exposures are smooth spatial functions, such as distance to pollution sources or major roadways, common in environmental studies.
@@ 1 Introduction
Estimating the coefficient corresponding to a structural linear spatial regression of a spatially varying outcome on a spatially varying exposure is a fundamental problem in spatial statistics and environmental epidemiology. However, this task is challenging in the presence of unmeasured spatial variables that simultaneously influence both the exposure (covariate) and the outcome. When such variables are not accounted for, the (conditional) effect size in standard regression models applied to point-referenced spatial data is unidentifiable, leading to biased estimates. This phenomenon is referred to as spatial confounding, and it arises frequently in real-world applications where spatial structure is intrinsic to both the exposures and to unmeasured background processes.
@@
An early formal recognition of this issue can be found in clayton1993spatial, which demonstrated how including spatially varying error terms in the model changes covariate effect estimates, and this change was favored over the unadjusted estimate which is biased under confounding. Subsequent works have had differing opinions about this (woodard1999estimating, reich2006effects, wakefield2007disease, hodges2010adding, schnell2020mitigating, khan2022restricted, zimmerman2022deconfounding) with arguments put forth both for and against including spatially correlated errors in the model. A particularly influential contribution by paciorek2010importance quantified the bias due to spatial confounding in generalized least squares (GLS) settings, relating it to the relative smoothness of the exposure and the unmeasured confounder. This work also underscored the identifiability challenges of separating the exposure effect from the influence of spatial confounding, unless certain restrictions or structural assumptions are imposed. Extensive theoretical and empirical results in khan2023re largely agree with the conclusions of paciorek2010importance that it is better to include spatial error terms in the analysis model when the exposure is rougher than the unmeasured confounder. page2017estimation arrives at similar conclusions for spatial predictions. gilbert2025consistency proved that the GLS estimator is consistent under varied forms of spatial confounding as long as the exposure has some non-spatial variation. dupont2022spatial+ proves consistency of their spatial+ method under a similar assumption of Gaussian noise in the exposure. gilbert2021causal grounded the spatial confounding problem in the framework of causal inference, going beyond the linear regression case. Their sufficient conditions for identifiability of a causal effect under confounding also assumed some non-spatial variation in the exposure.
The story is more complicated in settings where the exposure of interest is a smooth function of spatial location, with no non-spatial component. Estimators like GLS and spatial+, the consistency of which is guaranteed in the presence of noise in the exposure, may no longer work in this smooth-exposure setting. When exposures are too smooth relative to the confounder, the exposure effect cannot be identified (bolin2025spatial, datta2025consistent), but many methods make progress via the assumption that unmeasured confounders vary mainly at broad spatial scales, while exposures display more localized, high-frequency spatial patterns (thaden2018structural, keller2020selecting, dupont2022spatial+, dupont2023demystifying, guan2023spectral). In this case the exposure is unconfounded at higher frequencies and, along with the assumption of a linear relationship between the outcome and the exposure, common in spatial regression, this solves the problem of spatial confounding.
@@
When the unmeasured confounder has the same or finer spatial variation than the exposure, existing techniques generally fail, and standard spatial estimators can perform worse than even the unadjusted estimator as seen in paciorek2010importance, khan2023re. This can be viewed through the lens of causal inference as a violation of the positivity assumption, which requires that all levels of exposure be observed across levels of confounders (schnell2020mitigating). Examples of exposures that themselves vary at broad spatial scales are common in many environmental studies, e.g., distance to fracking mines (casey2016unconventional, currie2017hydraulic) or to roadways (karner2010near). brugge2013highway explicitly raises concerns about the inability to adjust for co-smooth spatial confounders. Furthermore, the notion of smoothness defined by an ordering of basis functions may not generalize to two or higher-dimensional spatial domains. For example, using the Kronecker product of basis functions of one-dimensional bases creates a two-dimensional basis with two indices dictating smoothness (one along each coordinate); Approaches like those of keller2020selecting, which model the confounder using the lowest indexed basis functions, do not easily generalize for such basis sets.
@@
In this work, we develop a general framework for effect estimation in spatial linear regression models in the presence of unmeasured spatial confounding, without relying on the assumption that confounding operates more smoothly in space than the exposure. As far as we are aware, ours is the first approach to spatial confounding to relax this assumption. We introduce a plurality-rule assumption inspired by ideas in the robust instrumental variable (IV) literature. The central assumption is that, in a representation of the exposure $X$ and the confounder $U$ along a set of orthonormal basis functions, there are enough basis functions that support the exposure but not the confounder. More precisely, we assume that the set of ratios of the projections of $U$ and $X$ onto each of the basis functions has a larger cluster around zero than around any other value. This assumption subsumes the typical assumption of finer scale variation in the exposure as a special case. Under the plurality assumption, we show that a set of candidate estimators of the regression effect, one for each basis function, will have the mode at the true coefficient value. We prove that this enables consistent estimation of the effect without prior knowledge of which basis functions support only the exposure. We call the method basis voting, as each basis yields a candidate estimate, and the most popular candidate – the mode – is the final estimate.
@@
A second contribution of this work is to improve the efficiency of the candidate estimators relative to projection-based estimators that have been used in the IV literature and in a plurality-rule-based deconfounding method recently proposed for time-series data (schur2025decor). We show that in our setting the projection-based candidate estimator is inefficient and present a novel drop-one candidate estimator with lower asymptotic variance and finite-sample unbiasedness. We prove consistency and asymptotic normality of the valid candidate estimators and consistency of the aggregate mode estimator. We empirically demonstrate robustness to the choice of basis functions used for analyzing the data.
@@ 2 Problem Setup and Assumptions
Let $\mathcal{D}\subseteq\mathbb{R}^{d}$ denote a compact spatial domain, and let $\bm{S}_{n}=\{S_{1},\dots,S_{n}\}$ represent $n$ spatial locations sampled from a distribution with density $f_{S}(\cdot)$ on $\mathcal{D}$. At each location $S_{i}$, we observe an exposure value $X_{i}\coloneqq X(S_{i})$ and an outcome value $Y_{i}\coloneqq Y(S_{i})$. We denote the exposure and outcome vectors by $\bm{X}_{n}=(X_{1},\dots,X_{n})^{\top}$ and $\bm{Y}_{n}=(Y_{1},\dots,Y_{n})^{\top}$, respectively. In addition to $X$ and $Y$, we assume there exists an unmeasured confounder $U$ defined at each location, $U_{i}\coloneqq U(S_{i})$, represented by the vector $\bm{U}_{n}=(U_{1},\dots,U_{n})^{\top}$, which is correlated with the exposure and affects the outcome. We assume a univariate exposure, and without loss of generality under the linear model the confounder can be represented by univariate $U$.
In keeping with the existing literature on spatial statistics and spatial confounding, we assume throughout a linear data generating process (DGP) for the outcome at a generic spatial location $s\in\mathcal{D}$:
@@
where $\beta$ is the exposure effect of interest, $\epsilon(s)$ is an independent mean-zero error term, and $U(s)$ is collinear with $X(s)$. If the confounder $U$ were measured, then it would be straightforward to estimate the exposure effect $\beta$, simply by regressing $Y$ in $X$ and $U$ and looking at the estimated slope for $X$. As $U$ is not measured, we cannot directly adjust for it. Both $X$ and $U$ are assumed to be deterministic (fixed) continuous functions of the spatial domain from a functional space $\mathcal{F}(\mathcal{D},f_{S})$ of continuous functions in $\mathcal{D}$ equipped with the inner product $\langle g,h\rangle_{f_{S}}\coloneqq\int_{\mathcal{D}}g(s)h(s)f_{S}(s)ds$ and the corresponding norm $\|\cdot\|_{f_{S}}$ with respect to the sampling measure $f_{S}$. Let $\mathcal{H}=\{h_{1},h_{2},\cdots\}$ denote an infinite-dimensional orthonormal basis of $\mathcal{F}(\mathcal{D},f_{S})$. Unlike prior approaches, our method does not require an assumption that the basis functions $h_{1},h_{2},\cdots$ are ordered by their degree of smoothness or roughness, with $X$ having a longer expansion than $U$.
Let $X(s)=\sum_{j=1}^{\infty}\alpha_{x}^{(j)}h_{j}(s)$ and $U(s)=\sum_{j=1}^{\infty}\alpha_{u}^{(j)}h_{j}(s)$. As they are continuous functions on a compact domain, they have finite $L^{2}$ norms: $\mathbb{E}[X(S)]^{2}=\sum_{j}\left|\alpha_{x}^{(j)}\right|^{2}<\infty$ and $\mathbb{E}[U(S)]^{2}=\sum_{j}\left|\alpha_{u}^{(j)}\right|^{2}<\infty$. By orthogonality, the projections of $X$ on each of the basis functions are given by the coefficients $\alpha_{x}^{(j)}=\langle X,h_{j}\rangle_{f_{S}}$ and similarly for $\alpha_{u}^{(j)}$.
Without loss of generality, we partition $\mathcal{H}$ into three disjoint sets $\mathcal{H}=\mathcal{H}_{UX}\sqcup\mathcal{H}_{X}\sqcup\mathcal{H}_{U}$ where $\mathcal{H}_{UX}$ is the set of basis functions shared by both $U$ and $X$, $\mathcal{H}_{X}$ is the set unique to $X$ and similarly for $\mathcal{H}_{U}$. Figure 1 depicts our DGP, with and without the basis expansion.
@@
(a) Standard confounding setting
$\mathcal{H}_{X}$$\mathcal{H}_{UX}$$\mathcal{H}_{U}$UUXXYY
(b) Expanded setting for spatial confounding with basis components (shaded grey)
@@
Assumption 1 (Oracle identification assumption).
Let $\bm{\alpha}_{x}=(\alpha_{x}^{(1)},\alpha_{x}^{(2)},\dots)$ and $\bm{\alpha}_{u}=(\alpha_{u}^{(1)},\alpha_{u}^{(2)},\dots)$ be the sequence of coefficients of $X$ and $U$ for the basis function set $\mathcal{H}$ respectively, then there exists a non-empty subset $O_{X}\subseteq\mathbb{N}$ such that for all $j\in O_{X}$, we have $\alpha_{u}^{(j)}=0$ and $\alpha_{x}^{(j)}\neq 0$.
If the index $j$ of even a single basis function $h_{j}\in O_{X}$ were known, that would suffice to identify $\beta$. This is because the effect of $h_{j}$ itself on $Y$ is unconfounded, and under linearity, its effect is equal to $\beta$. In practice, however, this will not be the case as $O_{X}$ is unknown and hence this is only an oracle identification assumption. Previous methods have instead relied on the assumption that the higher frequency basis functions are unconfounded. More formally, let $d$ denote the number of basis functions used in the analysis. Then when using a sufficiently large $d$, these methods essentially assume that the basis functions are ordered in terms of increasing roughness or frequency $h_{1}\lessapprox h_{2}\lessapprox\cdots\lessapprox h_{d}$, and there exists some $t\leq d\in\mathbb{N}$ such that $\{t+1,\cdots,\}\in O_{X}$. So, Assumption 1 is strictly weaker as, rather than requiring unconfounded high-frequency components, we only assume the existence of a non-trivial set $\{h_{j}:j\in O_{X}\}$. Figure 2 shows two scenarios amenable to our method, one in which the assumption of higher frequency variation in $X$ holds and one in which it is violated.
@@
A helpful way to interpret the subset of basis functions $\{h_{j}:j\in O_{X}\}$ is through the lens of instrumental variable (IV) methods in causal inference. In IV frameworks, one seeks an instrument $Z$ that satisfies two key properties: (i) it is correlated with the exposure $X$, and (ii) it is independent of the unmeasured confounder $U$ and the outcome $Y$, except through its effect on $X$. When such a valid instrument exists, projecting both $X$ and $Y$ onto the space spanned by $Z$ and regressing the projected outcome on the projected exposure yields a consistent estimate of the causal effect. In our spatial basis framework, the basis functions $\{h_{j}\}$ can be thought of as candidate instruments. Under Assumption 1, those $h_{j}$ with $j\in O_{X}$ are akin to valid instruments, in the sense that they influence $X$ but not $U$, and hence satisfy the classical exclusion restriction. The remaining basis functions $h_{j}$ for $j\notin O_{X}$ are invalid instruments, since they affect both $X$ and $U$, and their use would introduce bias into causal effect estimation.
To illustrate this connection, suppose we knew the set $O_{X}$, and let $O_{X,d}:=O_{X}\cap\{1,2,\ldots,d\}$ for a sufficiently large $d$ such that $O_{X,d}$ is non-empty. Then a least squares regression of $Y$ onto the basis functions in $O_{X,d}$, which gives a consistent estimator of $\beta$, is equivalent to the standard IV two-stage least squares procedure that projects $X$ onto the span of $\{h_{j}:j\in O_{X,d}\}$ and regresses $Y$ onto the predicted values.This equivalence is due to the fact that the projection of $X$ onto the span of $\{h_{j}:j\in O_{X,d}\}$ simply returns a linear combination of the basis functions in that set. This IV-based perspective is consistent with the framework of woodward2024instrumental, who show how a variety of existing approaches to unmeasured spatial confounding can be unified under a common IV interpretation.
@@
The challenge, of course, is that $O_{X}$ is unknown as we no longer assume that the unconfounded variation in $X$ is at the end of the frequency spectrum of the chosen set of $d$ basis functions. Since we do not observe $U$, we cannot directly verify which basis functions fall in the support of $X$ but not $U$. In the IV literature, several recent approaches have been developed to address this problem under a majority-rule (kang2016instrumental, windmeijer2019use) or plurality-rule (hartwig2017robust, guo2018confidence) assumption, namely that more than 50% of candidate instruments (majority rule) or that the largest subset of candidate instruments (plurality rule) result in estimators with the same unconfounded probability limit. A similar idea was proposed in the time-series setting with unmeasured temporal confounders by schur2025decor. They propose a DecoR estimator that projects both $X$ and $Y$ onto a collection of basis functions and defines aggregate estimators from these projections based on a majority-rule assumption. They frame the problem as an adversarial outlier problem and use the Torrent algorithm (bhatia2015robust) in a robust regression framework to estimate the effect size, based on a user-specified upper bound on cardinality of the set of invalid candidates, which is assumed to be small relative to the sample size. We compare DecoR to our method in a one-dimensional time series simulation setup in the supplementary material (S1) and see that DecoR fails to recover the effect when the majority rule does not hold but the plurality rule does.
@@
Inspired by these ideas, we propose a plurality-rule estimator which is defined as the mode of a set of candidate estimators, where each candidate is constructed based on each of the spatial basis functions $\{h_{j}\}_{j=1}^{d}$. This approach allows us to recover $\beta$ even without knowledge of $O_{X}$, under a plurality-rule assumption that enough of the candidate basis functions are valid, i.e., contribute to $X$ but not $U$. Most of the existing methods based on majority- or plurality-rule assumptions use a two-stage approach with sample splitting, in which they first select a set of valid candidates and then use them in a second estimation step. In contrast, we propose a mode-based estimator that obviates sample splitting as it is often hard to justify in spatial setting. Instead, our drop-one candidate estimator includes invalid basis functions as precision variables to increase efficiency. hartwig2017robust proposed a similar mode-based estimator with projection-based candidates in the Mendelian randomization IV setting, however they did not provide any theoretical guarantees for their procedure. In the next section, we formally define the plurality-rule assumption and our estimator, present its theoretical guarantees, and describe its advantages in comparison to existing approaches.